# Jingyao Liu♠♣  Chen Huang♠♢†  Zhizhao GuanWenqiang Lei♠♣Yang Deng

### v1

@@ B.2 Implementation of Evaluation Metrics
    where $\alpha$ and $\beta$ control the relative importance of requirement-level and test-level correctness, respectively. In this paper, we set $\alpha=0.6$ and $\beta=0.4$, reflecting our intuition that requirement-level accuracy (Req.Acc) plays a more critical role in evaluating the overall performance than test-level accuracy (Test.Acc).
@@
    $\text{Cost}=(\text{Input Tokens}\times\text{Input Price})+(\text{Output Tokens}\times\text{Output Price})$ (4)
@@
We randomly selected 10 tasks from the E2EDev benchmark and evaluated their implementations under 30 different models (5 backbones $\times$ 6 frameworks), resulting in 300 model-task instances. We recruited 5 professional annotators with expertise in software testing. To ensure labeling consistency, 5 shared examples were used to calibrate judgments, and the remaining 295 instances were evenly distributed across annotators.
@@
Table 7: Error types identified during human evaluation
@@ Appendix C Additional Analysis
C.1 Additional Effectiveness Analysis on E2EDev
@@
C.2 Additional Efficiency Analysis on E2ESD
@@
C.3 Additional Analysis on HumanEval (Function-level Task)
The experimental results presented in the main text indicate that current methods are not effective in handling project-level ESESD tasks. To gain further insights, this chapter shifts focus to evaluate the performance of these methods on simpler, function-level tasks, using the benchmark dataset HumanEval (Chen et al., 2021).
As summarized in Table 10, there is a noticeable improvement in the performance of various methods on the HumanEval benchmark compared to their results on the more complex E2ESD task. This suggests that current approaches are still primarily effective for simpler, function-level code generation tasks. It is worth noting that MetaGPT’s performance on HumanEval aligns closely with the results reported in its original paper (see the supplementary material of their work for details).
@@ C.3 Additional Analysis on HumanEval (Function-level Task)
C.4 Case Studies on LLM-driven Methods for E2ESD
@@ D.1 Test Case Specification – Gherkin
Scenario: User clicks the "I’m HUNGRY" button
@@
When the user clicks the "I’m HUNGRY" button
@@
Then the display should show the message "I’m HUNGRY"
Listing 1: Example Gherkin Test Case with Comments
@@ D.2 Test Script & Automated Testing Framework – Behave
Listing 2: Python Step Implementation for Gherkin Scenario
@@ E.3 Comparison
Automation Executed by frameworks like Behave or Cucumber, mapping natural-language scenarios to code steps Executed within testing frameworks like unittest or pytest
@@
Table 11: Comparison between BDD tests and Unit tests
@@ Appendix G Prompts Details
Requirement Extractor """system prompt""" You are an expert in extracting **functional** user requirements from web applications. Generate a **comprehensive and testable** list of user requirements that cover all user-facing functionalities. ### Functional Requirement Criteria Each requirement must include the following elements to ensure it is complete and testable: 1. **ID**: A unique identifier (e.g., REQ-001). 2. **Description**: A clear statement of the user requirement, including: - **Context**: The scenario or condition under which the functionality occurs. - **User Action**: What the user does (e.g., clicks, types, scrolls). - **System Response**: The expected outcome after the user action. ### **Rules** - Only include **functional requirements** - i.e., observable behaviors triggered by user interaction via the front end (such as clicking a button, entering text, receiving visual feedback, etc). - **Avoid including non-functional requirements** such as performance, security, or scalability unless they are visible or interactive on the UI. - Exclude backend logic unless it has a **direct effect on the UI** that is visible or interactive. ### **Output Format (JSON)** { "summary": { "overview": "Briefly describe the applications purpose and key functionalities.", "predefined_options": "Predefined options set by the system to standardize inputs and reduce manual configuration, such as default values and preset selections.", "external_resources": "External resources used by the application, including links, images, audio files, and other media. List the resource names and their sources (URLs or file paths).", "external_js_libraries": "External JavaScript libraries or packages used by the application, such as jQuery, React, Bootstrap, etc. Provide the library names and their sources (e.g., CDN links)." } , "requirements": [ { "id": "Unique identifier", "description": "User requirement description with context, user action, and system response." } ] }

### v4

@@ B.2 Implementation of Evaluation Metrics
    where $\alpha$ and $\beta$ control the relative importance of requirement-level and test-level correctness, respectively. In this paper, we set $\alpha=0.6$ and $\beta=0.4$, reflecting our intuition that requirement-level accuracy (Req.Acc) plays a more critical role in evaluating the overall performance than test-level accuracy (Test.Acc). Due to the limited number of benchmark tasks at the project level, we run each framework three times on every task. The effectiveness scores reported in Table 3 and Table 8 are computed as the mean and standard deviation across runs.
@@
    $\begin{split}\text{Cost}&=(\text{Input Tokens}\times\text{Input Price})\\ &\quad+(\text{Output Tokens}\times\text{Output Price})\end{split}$ (4)
@@
We randomly selected 10 tasks from the E2EDev benchmark and evaluated their implementations under 30 different models (6 backbones $\times$ 6 frameworks), resulting in 360 model-task instances. We recruited 5 professional annotators with expertise in software testing. To ensure labeling consistency, 5 shared examples were used to calibrate judgments, and the remaining 295 instances were evenly distributed across annotators.
@@
Table 7: Error types identified during human evaluation.
@@ Appendix C Additional Analysis
C.1 OpenHands and SWE-Agent on E2EDev
Backbone Method Effectiveness ($\uparrow$, %) Efficiency ($\downarrow$)
Req. Acc. Test Acc Bal. Score Cost (\$) CO2 (g) Time (s)
Claude-Haiku 4.5 Vanilla LLM 48.69 $\pm$1.30 63.08 $\pm$0.50 54.45 $\pm$0.98 0.015 0.028 53
GPT-Engineer 53.75 $\pm$0.43 69.41 $\pm$1.07 60.01 $\pm$0.69 0.016 0.038 59
Self-Collab. 49.01 $\pm$1.74 61.50 $\pm$1.26 54.01 $\pm$1.55 0.014 0.033 75
MapCoder 49.61 $\pm$0.50 65.65 $\pm$0.44 56.03 $\pm$0.48 0.099 0.081 129
ChatDev 44.73 $\pm$1.50 58.09 $\pm$1.03 50.07 $\pm$1.31 0.174 0.188 335
MetaGPT 5.39 $\pm$0.76 10.66 $\pm$1.77 7.50 $\pm$1.16 0.090 0.091 366
Openhands 43.45 57.77 49.18 0.021 0.054 -
SweAgent 44.73 $\pm$0.44 56.87 $\pm$0.45 49.59 $\pm$0.44 0.089 0.089 148
GPT-4o Vanilla LLM 45.95 $\pm$1.25 60.88 $\pm$1.97 51.92 $\pm$1.54 0.0160 0.083 28
GPT-Engineer 50.83 $\pm$1.33 66.59 $\pm$2.43 57.13 $\pm$1.77 0.0198 0.132 21
Self-Collab. 46.83 $\pm$1.12 61.15 $\pm$1.47 52.56 $\pm$1.26 0.0155 0.109 37
MapCoder 47.70 $\pm$2.57 63.97 $\pm$2.25 54.21 $\pm$2.44 0.1091 0.750 93
ChatDev 42.71 $\pm$2.44 58.93 $\pm$2.66 49.20 $\pm$2.53 0.1947 1.910 114
MetaGPT 0.00 $\pm$0.00 0.17 $\pm$0.05 0.07 $\pm$0.02 0.0951 0.794 66
Openhands 39.25 55.34 45.85 0.062 0.161 -
SweAgent 40.83 $\pm$2.23 58.05 $\pm$1.91 47.72 $\pm$2.10 0.089 0.089 122
Table 8: Benchmark analysis of effectiveness and efficiency. The grey row denotes the Vanilla baseline. Red and green indicate performance better and worse than Vanilla, respectively. Standard deviations are shown in small grey text. For effectiveness, the best result is shown in bold and the second-best is underlined.
Method Task Description Input Output Mode
E2EDev (Ours) Generates fully runnable software projects from scratch based solely on requirements. Detailed user requirements Runnable project End-to-End Repo Generation
OpenHands Collaborates via multi-turn dialogue to iteratively build software. Interactive dialogue Runnable project Human-in-the-loop Repo Generation
SWE-Agent Resolves GitHub issues by modifying existing repositories (SWE-bench). Existing repo + Issue Patched repo Repo Modification
Table 9: Comparison of task formulations. E2EDev targets ab initio generation, distinct from the collaborative or maintenance-focused nature of OpenHands and SWE-Agent.
OpenHands (Wang et al., 2024) and SWE-Agent (Yang et al., 2024) are two representative and widely used LLM-based agents for software engineering tasks. Both systems are designed to support complex development workflows and have shown strong performance in their original evaluation settings. However, their design assumptions and task formulations differ substantially from the setting targeted by E2EDev, as summarized in Table 9.
E2EDev focuses on end-to-end requirement-to-executable software development, where a system is required to generate a complete and runnable software project directly from natural language requirements. This setting emphasizes project-level generation from scratch. In contrast, OpenHands is built around iterative human–agent collaboration, where continuous human feedback guides the development process, while SWE-Agent operates on an existing code repository and targets issue-level modification rather than full project generation. As a result, these systems are not natively aligned with the E2EDev task formulation, and a direct comparison with E2EDev-oriented frameworks is not conceptually equivalent.
Experimental Setup.
Due to their substantially higher evaluation cost and different design goals, we evaluate OpenHands and SWE-Agent only on GPT-4o and Claude-Haiku-4.5. To better align these systems with the E2EDev setting, OpenHands is restricted to a single execution without human interaction. For SWE-Agent, we replace its original prompt template—which instructs the model to load and modify an existing repository—with a new template that directs the model to generate a runnable software project directly from the given user requirement. This modification prevents access to any existing repositories and enforces project generation from scratch.
Results and Discussion.
The results in Table 8 show that OpenHands and SWE-Agent underperform the Vanilla LLM baseline on E2EDev. This outcome is expected given the mismatch between their original design objectives and the requirement-to-executable software development task. In particular, both systems exhibit limitations in directly interpreting and implementing fine-grained user requirements without relying on interactive guidance or pre-existing code structures. Overall, these findings further support our main conclusion that E2EDev poses distinct challenges beyond those addressed by existing agent-based software engineering frameworks.
C.2 Additional Effectiveness Analysis on E2EDev
@@
claude-haiku 4.5 ChatDev 0.4464 0.7025 0.2718
gpt-engineer 0.5402 0.7991 0.2589
llmbased 0.4866 0.7454 0.2589
MapCoder 0.4955 0.7723 0.2768
MetaGPT 0.0536 0.1027 0.0491
Self-collaboration-Code-Generation 0.4911 0.7679 0.2768
@@
C.3 Additional Efficiency Analysis on E2ESD
@@
C.4 Additional Analysis on HumanEval (Function-level Task)
The main experimental results show that existing methods struggle to handle project-level E2ESD tasks. To better understand this limitation, we further evaluate these methods on a simpler, function-level benchmark, HumanEval (Chen et al., 2021), which isolates code generation from project-level coordination and integration.
As summarized in Table 12, most methods achieve substantially higher performance on HumanEval than on E2ESD. This gap indicates that current approaches remain more effective for isolated function-level code generation than for end-to-end project development.
For MetaGPT, we deliberately evaluate the framework without any HumanEval-specific adaptation. Under this setting, MetaGPT’s performance is lower than the results reported in the main body of its original paper, but closely matches the authors’ appendix results obtained without benchmark-specific tuning. This confirms that the observed performance is consistent with MetaGPT’s default framework design. A closer inspection reveals that most performance degradation arises from the multi-agent collaboration process rather than from code generation itself. During requirement analysis and architectural planning, MetaGPT frequently modifies the target function name or signature specified by the benchmark. As a result, the generated code no longer matches the expected test interface, causing test failures even when the core logic is otherwise correct. This behavior mirrors the failure patterns observed on E2EDev, where the framework’s internal analysis stages often alter the original requirements, leading to mismatches between the specified requirements and the generated implementation.
@@ C.4 Additional Analysis on HumanEval (Function-level Task)
C.5 Case Studies on LLM-driven Methods for E2ESD
@@ D.1 Test Case Specification – Gherkin
Listing 1: Example Gherkin Test Case with Comments
@@
Scenario: User clicks the "Im HUNGRY" button
@@
When the user clicks the "Im HUNGRY" button
@@
Then the display should show the message "Im HUNGRY"
@@ D.2 Test Script & Automated Testing Framework – Behave
Listing 2: Python Step Implementation for Gherkin Scenario
@@ E.3 Comparison
Automation Executed by frameworks like Behave, mapping scenarios to code Executed within testing frameworks like unittest or pytest
@@
Table 13: Comparison between BDD tests and Unit tests.
@@ Appendix G Prompts Details
Requirement Extractor """system prompt""" You are an expert in extracting **functional** user requirements from web applications. Generate a **comprehensive and testable** list of user requirements that cover all user-facing functionalities. ### Functional Requirement Criteria Each requirement must include the following elements to ensure it is complete and testable: 1. **ID**: A unique identifier (e.g., REQ-001). 2. **Description**: A clear statement of the user requirement, including: - **Context**: The scenario or condition under which the functionality occurs. - **User Action**: What the user does (e.g., clicks, types, scrolls). - **System Response**: The expected outcome after the user action. ### **Rules**: - Only include **functional requirements** - i.e., observable behaviors triggered by user interaction via the front end (such as clicking a button, entering text, receiving visual feedback, etc.). - **Avoid including non-functional requirements** such as performance, security, or scalability unless they are visible or interactive on the UI. - Exclude backend logic unless it has a **direct effect on the UI** that is visible or interactive. ### **Output Format (JSON)** { "summary": { "overview": "Briefly describe the applications purpose and key functionalities.", "predefined_options": "Predefined options set by the system to standardize inputs and reduce manual configuration, such as default values and preset selections.", "external_resources": "External resources used by the application, including links, images, audio files, and other media. List the resource names and their sources (URLs or file paths).", "external_js_libraries": "External JavaScript libraries or packages used by the application, such as jQuery, React, Bootstrap, etc. Provide the library names and their sources (e.g., CDN links)." } , "requirements": [ { "id": "Unique identifier", "description": "User requirement description with context, user action, and system response." } ] }